[figure: 9.7 cm]

POINT CONTACT SPECTROSCOPY OF THE MELT-SPUN

CUBIC ErCu_{5}

S. Ilkovič^{a}, M. Reiffers^{b}, B. Idzikowski^{c}, K.H. Müller^{d}

aFaculty of Sciences, University of Prešov, 17.novembra, SK-080 78 Prešov, Slovakia

bInstitute of Experimental Physics, Watsonova 47, SK-043 53 Košice, Slovakia cInstitute of Molecular Physics, Polish Academy of Sciences, M.Smoluchowskiego 17,

PL 60-179 Poznaň, Poland

dInstute für Festkörper Werkstofforschung Dresden, D-01171 Dresden, Germany

The intermetallic compound ErCu_{5} crystallizes in cubic structure. Up to now, there exist only a few data about ErCu_{5}. The probable existence of a ferromagnetic ordering at low temperatures below 2 K follows from the low temperature behaviour of magnetization. One the other hand one can supposes the influence of crystalline electric field (CEF) effect on the transport properties. Therefore we have performed the study of the electron-quasiparticle interaction function by point contact spectroscopy at low temperatures. The pollycrystalline samples have been prepared in ribbon shape by low temperature melt-spun and subsequent annealing. We present the point contact spectra of DyCu_{5} - Cu heterocontacts which are directly proportional to the electronquasiparticle interaction function. We have observed the main peak at about 19 meV which is characteristic for electron-phonon interaction function of other RECu_{5} (RE - heavy rare earths) compounds. Moreover, we have observed the additional maximum at about 2.5 meV which is connected probably with the allowed transition from CEF ground state to higher excited CEF level.

[figure: 13.4 cm]

Subject category :

4. Rare Earths and Actinides, Alloys and Compounds

Presentation mode : poster

Corresponding author : M. Reiffers

Address for correspondence : Institute of Experimental Physics Watsonova 47 SK-043 53 Košice Slovakia

Email address : reiffers@kosice.upjs.sk